### v1

@@ 6 Conclusion
In our doubly robust Granger causality test, the types of deep neural networks used are relatively basic. More complex architectures may yield improved results. For instance, increasing the depth of the network or adopting more advanced neural network models could potentially enhance the test’s performance. Additionally, our parameter settings for the neural networks are relatively flexible. When applying the method to different causality testing problems, experienced practitioners might be able to fine-tune the model to obtain more insightful conclusions. In these respects, our work may open up potential directions for future research.
@@
This appendix first summarizes the notation used throughout the appendix. Appendices A-C then contain the assumptions, technique lemmas, main proofs, and a comparison with the DGPs in Nishiyama et al. (2011), respectively.
We summarize the notation used in the appendix as follows:
  • We use $C,C_{1},C_{2},C_{m},C^{\prime},C^{\prime\prime},C_{\ell},c_{1},c_{2},\omega_{1},\omega_{2},\kappa,\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{B},r_{0},\rho_{0},\widetilde{\gamma},\eta$ to denote generic constants, whose values may vary from place to place. Let $B_{n}\asymp n^{\kappa_{B}}$ for some $\kappa_{B}\in(0,1/2)$.
  • The smoothness parameters are denoted by $\beta_{0}$ and $\gamma_{0}$. For parameters of the network, denote $W$ as the total number of parameters, $L_{n}$ as the depths (number of layers), and $H_{n}$ as the common asymptotic order of width that all layers share.
  • For the $g$th component of the conditional mixture Gaussian model in the MDN, the weight, mean, and variance parameters are written as $\alpha_{g}(\cdot),\mu_{g}(\cdot)$, and $\sigma_{g}(\cdot)$, respectively.
  • The loss function for training a deep neural network is denoted by $\ell(f,\boldsymbol{Z}_{t})$, where $f$ is the function input to the loss and $\boldsymbol{Z}_{t}=(Y_{t},\boldsymbol{Y}_{t-1}^{\top})^{\top}$ collects the response $Y_{t}$ and the covariates $\boldsymbol{Y}_{t-1}$. Define the true conditional expectation function $m(\boldsymbol{Y}_{t-1})=\operatorname*{arg\,min}\mathbb{E}[\ell(f,\boldsymbol{Z}_{t})]$. Denote $\widehat{m}(\boldsymbol{Y}_{t-1})$ as the MLP estimator.
  • Denote the true conditional density function of $\boldsymbol{X}_{t-1}$ given $\boldsymbol{Y}_{t-1}$ as $f_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(\cdot|\cdot)$. Let $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(\cdot|\cdot)$ be the conditional density estimated by the MDN. The corresponding estimated characteristic function is $\widehat{\phi}(\nu|\boldsymbol{Y}_{t-1})$ .
  • The function classes of deep neural networks and multilayer perceptrons are denoted by $\mathcal{F}_{\mathrm{DNN}}$ and $\mathcal{F}_{\mathrm{MLP}}$, respectively. See Anthony and Bartlett (1999), Goodfellow et al. (2016) and Farrell et al. (2021) for details.
  • Define
    $\widehat{f}_{\mathrm{DNN}}=\underset{\begin{subarray}{c}f\in\mathcal{F}_{\mathrm{DNN}}\\ \left\|f\right\|_{\infty}\leq 2C_{m}\end{subarray}}{\arg\min}\sum_{t=q+1}^{n}\ell\left(f,\boldsymbol{Z}_{t}\right).$ (13)
[…]
    $f_{n}=\underset{\begin{subarray}{c}f\in\mathcal{F}_{\mathrm{DNN}}\\ \|f\|_{\infty}\leq 2M\end{subarray}}{\arg\min}\left\|f-m\right\|_{\infty},$
    and let $\epsilon_{n}=\left\|f_{n}-m\right\|_{\infty}$.
  • For any function $f(\cdot)$, we write $\|f(\cdot)\|_{\infty}$ for its sup-norm. For any vector $\boldsymbol{v}$, denote $\|\boldsymbol{v}\|_{1}$ for its $l_{1}$-norm.
@@ Appendix A Assumptions
Assume $X_{t}$ and $Y_{t}$ are stationary, and satisfy $\beta$-mixing condition, where the mixing coefficient $\beta(t)\leq C_{1}\exp\left(-C_{2}t\right)$ for some constants $C_{1},C_{2}>0$. Denote $\varepsilon_{t}^{*}=Y_{t}-\mathbb{E}(Y_{t}|\boldsymbol{W}_{t-1})$. Suppose $\{\varepsilon_{t}^{*},\mathcal{F}_{t}\}$ is a martingale difference sequence.
@@
Appendix B Technical lemmas
@@ B.2 Lemma 2
Suppose Assumptions 1, 3 and 4 hold. Suppose the MLP is specified with width $H_{n}\asymp n^{\left(\frac{q}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{2}(n)$ and depth $L_{n}\asymp\log(n)$. Let $\epsilon_{n}=n^{-\left(\frac{\beta_{0}}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{6}(n)$. Then, for the MLP estimator $\widehat{m}(\boldsymbol{Y}_{t-1})$, there exist constants $C,C_{1},C_{2}>0$ independent of $n$, for all $n$ sufficiently large, such that
$P\left(\mathbb{E}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}\epsilon_{n}^{2}\right)\geq 1-e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{2C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n),$
$P\left(\mathbb{E}_{n}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}\epsilon_{n}^{2}\right)\geq 1-4e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{12C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n).$
@@ B.3 Lemma 3
Suppose Assumptions 1 and 5 hold. Then, there exists a certain MDN function class satisfying Assumption 6, such that the resulting MDN estimator $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$ satisfies that, with probability at least $1-O\left(n^{-1}\right)$,
@@
for some constants $C,\omega_{1},\omega_{2}>0$.
@@ B.4 Lemma 4
Suppose Assumption 1 and 7 (ii)-(iv) hold. Then, there exists a constant $\kappa>1/2$, such that
@@ C.2 Proof for Proposition 2
Suppose Assumption 1, 7 and 8 hold. Then, under the null hypothesis,
@@ C.3 Proof for Theorem 1
Suppose Assumption 1, 7 and 8 hold. Then under the null,
@@ C.4 Proof for Corollary 1
Suppose Assumption 1, 7 and 8 hold. Then under the null,
@@ C.5 Proof for Theorem 2
Suppose Assumption 1, 7 and 8 hold. Then under the alternative, for each $(\mu,\nu)\in\mathcal{W}$,
@@ C.6 Proof for Theorem 3
Suppose Assumption 1, 7, 8 and 9 hold. Then under the sequence of local alternatives,
@@ C.7 Proof for Corollary 2
Suppose Assumption Assumption 1, 7, 8 and 9 hold. Then, under the sequence of local alternatives,
@@ C.8 Proof for Proposition 3
Suppose Assumption 1, 7 and 8 hold. The multiplier bootstrap empirical process satisfies
@@ C.9 Proof for Theorem 4
Suppose Assumption 1, 7 and 8 hold. Then, under the null, under the alternative, or under the sequence of local alternatives,

### v2

@@ 6 Conclusion
In our doubly robust Granger causality test, the types of deep neural networks used are relatively basic. More complex architectures may yield improved results, as in Zhang et al. (2025). For instance, increasing the depth of the network or adopting more advanced neural network models could potentially enhance the test’s performance. Additionally, our parameter settings for the neural networks are relatively flexible. When applying the method to different causality testing problems, experienced practitioners might be able to fine-tune the model to obtain more insightful conclusions. In these respects, our work may open up potential directions for future research.
@@
This appendix first summarizes the assumptions in Appendix A. Appendices B and C then contain the technique lemmas, main proofs, respectively.
@@ Appendix A Assumptions
Assume $X_{t}$ and $Y_{t}$ are stationary, and satisfy $\beta$-mixing condition, where the mixing coefficient $\beta(t)\leq C_{1}\exp\left(-C_{2}t\right)$ for some constants $C_{1},C_{2}>0$. Suppose $\{Y_{t}-\mathbb{E}(Y_{t}|\boldsymbol{W}_{t-1}),\mathcal{F}_{t}\}$ is a martingale difference sequence.
@@
Appendix B Proofs of lemmas
@@ B.2 Lemma 2
Suppose Assumptions 1, 3 and 4 hold. Suppose the MLP is specified with width $H_{n}\asymp n^{\left(\frac{q}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{2}(n)$ and depth $L_{n}\asymp\log(n)$. Let $K_{n}=n^{-\left(\frac{\beta_{0}}{\beta_{0}+q}\right)\left(1/2-\kappa_{B}\right)}\log^{6}(n)$. Then, for the MLP estimator $\widehat{m}(\boldsymbol{Y}_{t-1})$, there exist constants $C,C_{1},C_{2}>0$ independent of $n$, for all $n$ sufficiently large, such that
$P\left(\mathbb{E}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}K_{n}^{2}\right)\geq 1-e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{2C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n),$
$P\left(\mathbb{E}_{n}\left|\widehat{m}(\boldsymbol{Y}_{t-1})-m(\boldsymbol{Y}_{t-1})\right|^{2}\leq C^{2}K_{n}^{2}\right)\geq 1-4e^{-n\left(\frac{\beta_{0}}{\beta_{0}+q}\right)^{\left(1/2-\kappa_{B}\right)}}-\frac{12C_{1}n^{1-C_{2}\log(n)}}{\log(n)}-o(1/n).$
@@ B.3 Lemma 3
Suppose Assumptions 1 and 5 hold. Then, there exists a certain MDN function class satisfying Assumption 6, such that the resulting MDN estimator $\widehat{f}_{\boldsymbol{X}_{t-1}|\boldsymbol{Y}_{t-1}}(x|y)$ satisfies that
@@
for some constants $C,\omega_{1},\omega_{2}>0$, with probability at least $1-O\left(n^{-1}\right)$.
@@ B.4 Lemma 4
Suppose Assumptions 1 and 7 (ii)-(iv) hold. Then, there exists a constant $\kappa>1/2$, such that
@@ C.2 Proof for Proposition 2
Suppose Assumptions 1, 7 and 8 hold. Then, under the null hypothesis,
@@ C.3 Proof for Theorem 1
Suppose Assumptions 1, 7 and 8 hold. Then under the null,
@@ C.4 Proof for Corollary 1
Suppose Assumptions 1, 7 and 8 hold. Then under the null,
@@ C.5 Proof for Theorem 2
Suppose Assumptions 1, 7 and 8 hold. Then under the alternative, for each $(\mu,\nu)\in\mathcal{W}$,
@@ C.6 Proof for Theorem 3
Suppose Assumptions 1, 7, 8 and 9 hold. Then under the sequence of local alternatives,
@@ C.7 Proof for Corollary 2
Suppose Assumptions 1, 7, 8 and 9 hold. Then, under the sequence of local alternatives,
@@ C.8 Proof for Proposition 3
Suppose Assumptions 1, 7 and 8 hold. The multiplier bootstrap empirical process satisfies
@@ C.9 Proof for Theorem 4
Suppose Assumptions 1, 7 and 8 hold. Then, under the null, under the alternative, or under the sequence of local alternatives,